### v1

# MicroDreamer: Zero-shot 3D Generation in $\sim$20 Seconds by Score-based Iterative Reconstruction

Abstract
*Equal Contribution.
Corresponding author.
Optimization-based approaches, such as score distillation sampling (SDS), show promise in zero-shot 3D generation but suffer from low efficiency, primarily due to the high number of function evaluations (NFEs) required for each sample. In this paper, we introduce score-based iterative reconstruction (SIR), an efficient and general algorithm for 3D generation with a multi-view score-based diffusion model. Given the images produced by the diffusion model, SIR reduces NFEs by repeatedly optimizing 3D parameters, unlike the single optimization in SDS, mimicking the 3D reconstruction process. With other improvements including optimization in the pixel space, we present an efficient approach called MicroDreamer that generally applies to various 3D representations and 3D generation tasks. In particular, retaining a comparable performance, MicroDreamer is 5-20 times faster than SDS in generating neural radiance field and takes about 20 seconds to generate meshes from 3D Gaussian splitting on a single A100 GPU, halving the time of the fastest zero-shot baseline, DreamGaussian. Our code is available at https://github.com/ML-GSAI/MicroDreamer.
Keywords:
3D Generation Diffusion Model Multi-view Diffusion Score Distillation Sampling Zero-shot
Figure 1: MicroDreamer can generate high-quality meshes in about 20 seconds on a single A100, built on a multi-view diffusion model without additional 3D data.
1 Introduction
Recently, optimization-based approaches [40, 56, 19, 66, 5, 24, 46, 60, 55, 18, 53, 33, 62] particularly score distillation sampling (SDS) [40, 56] have emerged as promising avenues for 3D generation based on text-to-image diffusion models [48, 11, 51, 43, 44, 42]. These approaches are appealing due to their minimal, or even zero reliance on 3D data, in contrast to the data-intensive requirements of feed-forward approaches [3, 4, 6, 14, 37, 26, 13, 35, 59, 61, 65, 54, 23, 22, 67]. This advantage is particularly significant given that 3D data are costly and scarce.
Despite their promising capabilities, optimization-based approaches suffer from low efficiency. Intrinsically, the inefficiency lies in the extensive number of function evaluations (NFEs), i.e. forward passes of the diffusion model, required for each 3D object generation. Besides, within the commonly employed latent diffusion model (LDM) [43], the gradients of the 3D parameters need to be backpropagated through an encoder that maps rendered images to the latent space, further exacerbating the inefficiency of these approaches.
To address the computational issues, this paper introduces an efficient and general algorithm termed score-based iterative reconstruction (SIR) for zero-shot 3D generation in Sec. 3. Similar to SDS, SIR iteratively updates 3D parameters, leveraging a multi-view diffusion model without additional 3D data. However, in each iteration, SIR distinguishes itself by repeatedly optimizing 3D parameters given a single set of images produced by the diffusion model, mimicking the 3D reconstruction process to reduce the total NFEs (see Fig. 3(a)). To obtain 3D-consistent and high-quality images as the ground truth for 3D reconstruction in each iteration, we carefully design a hybrid forward process and a sampling process to refine the images rendered from the 3D object optimized in the latest iteration. Besides, even mapped back to the pixel space through the decoder in LDM, the refined images are still of high quality for reconstruction, enabling optimization in the pixel space to further speed up (see Fig. 6(a)-6(b)).
We delve into critical components beyond the algorithm to ensure compatibility with SIR. The comprehensive system is named by MicroDreamer, highlighting its efficiency for 3D generation. As detailed in Sec. 4, we introduce an initialization strategy for 3D objects, an annealed time schedule for diffusion, additional losses on reference images for image-to-3D, and a refinement procedure for deriving high-quality meshes from 3D Gaussian splatting (3DGS)[15].
SIR and MicroDreamer are broadly applicable to both neural radiance fields (NeRF)[34] and 3DGS, and to both text-to-3D and image-to-3D tasks, as detailed in Sec. 5. Employing three widely adopted multi-view diffusion models [47, 52, 58], we systematically compare SIR and SDS for NeRF generation. Retaining a comparable performance, SIR can accelerate the generation process by 5 to 20 times. Besides, MicroDreamer can efficiently produce 3DGS and further refine it into high-quality meshes, delivering consistent 3D meshes in about 20 seconds on a single A100 GPU – about twice as fast as the most competitive optimization-based alternatives, DreamGaussian [55]. Remarkably, MicroDreamer is on par with speed compared to feed-forward methods trained on extensive 3D data, with a very competitive CLIP similarity [41, 55].
Figure 2: Overview of SIR and SDS. $\mathcal{E}$, $\mathcal{D}$ and $\epsilon_{\phi}$ are the encoder, decoder, and diffusion model in LDM [43]. The solid and dashed arrows represent forward and backward passes respectively. The computation bottlenecks are highlighted in red. SIR is efficient because it lowers the NFEs in total and enables optimization in the pixel space.
2 Background
We present background on 3D representations, diffusion models, multi-view diffusion models, and current optimization-based 3D generation methods.
2.1 3D Representation
Neural radiance fields[34, 36] (NeRF) and 3D Gaussian splatting[15] (3DGS) have emerged as popular 3D representations. NeRF employs an MLP to predict the colors and density of the input space coordinates. 3DGS consists of multiple 3D Gaussians parameterized by the colors, center, scale, and rotation quaternion. We denote the corresponding tunable parameters in both representations as $\theta$. Given camera poses $c$, both approaches define a differential rendering process, denoted by $g(\theta,c)$. They are proven efficient and effective in 3D reconstruction [34, 15] and generation [40, 55].
2.2 Diffusion Models
A diffusion model [48, 11, 51] consists of a forward process and a sampling process. The forward process gradually adds Gaussian noise to an input image from time 00 to $T$. For any $t\in(0,T)$, the noise-adding process can be written as follows:
$x_{t}=\alpha_{t}x_{0}+\sigma_{t}\epsilon:=\textrm{Noise-adding}(x_{0},0\rightarrow t),\quad\epsilon\in\mathcal{N}(0,I),$ (1)
where the coefficients $\alpha_{t}$ and $\sigma_{t}$ form the noise schedule of the diffusion model. A noise prediction network $\epsilon_{\phi}(x_{t},t)$ with parameters $\phi$ is trained to predict the noise in the input $x_{t}$ with the corresponding noise level, i.e. time $t$. Plugging in the noise prediction network into Eq. (1), we can solve $x_{0}$ from a noisy image $x_{t}$ of time $t$ by a single-step prediction as follows:
$\hat{x}^{t}_{0}=\frac{1}{{\alpha}_{t}}x_{t}-{\frac{\sigma_{t}}{{\alpha}_{t}}}\epsilon_{\phi}(x_{t},t),$ (2)
which is an efficient approximation of the original input.
Instead of using Eq. (2) directly, the sampling process of diffusion gradually denoises through the noise prediction network and generates images from pure Gaussian noise. Among existing samplers [21, 2, 1, 30, 31, 64], the denoising diffusion implicit models (DDIM) [49] facilitate a sequence of samplers with random noise control $\eta$. When $\eta=0$, it solves the equivalent probability flow ordinary differential equation (ODE) [51] of the diffusion model and enjoys a fast sampling process with fewer number of function evaluations (NFEs) [^1] Throughout the paper, we refer to the number of forward passes through $\epsilon_{\phi}$ as NFEs.. In this setting, the one-step sampling of DDIM is given by:
$x_{t-1}=\frac{\alpha_{t-1}}{\alpha_{t}}x_{t}+\left(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t}\right)\epsilon_{\phi}(x_{t},t):={\textrm{Sampler}}(x_{t},t\rightarrow t-1).$ (3)
We refer the readers to the original paper [49] for the formula of $\eta>0$. For $\eta=1$, it represents a standard SDE sampler [11], with a higher tolerance for mismatches in the distributions of latent variables [29, 39].
Besides, when $\eta=0$ there is an inverse process (called DDIM inversion) that maps the distribution of images to the same distributions of noisy images as in Eq. (1) but it can maintain the unique feature of an input image and reconstruct it accurately through the corresponding DDIM sampler in Eq. (3). The (one-step) DDIM inversion [49] is formulated as follows:
$x_{t+1}=\frac{\alpha_{t+1}}{\alpha_{t}}x_{t}+\left(\sigma_{t+1}-\frac{\alpha_{t+1}}{\alpha_{t}}\sigma_{t}\right)\epsilon_{\phi}(x_{t},t):={\textrm{Inversion}}(x_{t},t\rightarrow t+1).$ (4)
2.3 Multi-view Diffusion Models
After being trained on a modest amount of 3D data, diffusion models can generate 3D-consistent multi-view, known as multi-view diffusion models. Among them, MVDream [47] takes text inputs and outputs multi-view images consistent in 3D. In contrast, Zero-1-to-3 [24] and ImageDream [58] focus on the Image-to-3D task, taking an additional reference image and its viewpoint as input. These models output new viewpoint images consistent with the reference image. This paper directly utilizes these pre-trained multi-view diffusion models without any further fine-tuning. For more details, please refer to the original papers and the source code in our supplementary materials. Notably, such multi-view diffusion models cannot provide a sufficient number of consistent multi-view images for classical 3D reconstruction, as shown in Fig. 6(c).
2.4 Optimization-based Algorithms for 3D Generation
Built upon (multi-view) diffusion models, optimization-based methods (see Sec. 6 for a review) aim to generate 3D content in a zero-shot manner. Among them, score distillation sampling (SDS) [40, 56] is the most representative and popular approach. Formally, denoting the rendered images as $x=g(\theta,c)$, SDS repeats adding noise to $x$ according to Eq. (1) and updating the 3D parameters $\theta$ by
$\nabla_{\theta}\mathcal{J}_{\textrm{SDS}}(x=g(\theta);\phi):=\mathbb{E}_{t}\left[w(t)(\epsilon_{\phi}(\alpha_{t}x+\sigma_{t}\epsilon,t)-\epsilon)\frac{\partial x}{\partial\theta}\right],$ (5)
where $w(t)$ is a fixed weighting function and we omit the dependency on the prompt $y$ and camera $c$ for simplicity. Notably, by reparameterization [66] according to Eq. (2), SDS has an equivalent data-prediction form:
$\nabla_{\theta}\mathcal{J}_{\textrm{SDS}}(x=g(\theta);\phi)=\mathbb{E}_{t}\left[\frac{w(t)\alpha_{t}}{\sigma_{t}}(x-\hat{x}_{0}^{t})\frac{\partial x}{\partial\theta}\right].$ (6)
3 Score-based Iterative Reconstruction
We present the general framework of score-based iterative reconstruction (SIR), an efficient and versatile algorithm inspired by classical 3D reconstruction in Sec. 3.1. SIR is designed to minimize the number of function evaluations (NFEs) typically required by existing methods. In Sec. 3.2, we introduce a diffusion-based process to produce multi-view images as ground truth for reconstruction. In Sec. 3.3, we demonstrate how SIR facilitates the optimization of 3D content directly within pixel space, thereby further boosting efficiency.
3.1 Mimicking 3D Reconstruction to Reduce the NFEs
Existing zero-shot 3D generation algorithms [40, 56, 66, 60] iteratively optimize the 3D parameters based on a 2D diffusion (see Eq. (5) for SDS as an example). Consequently, these methods necessitate a high number of function evaluations (NFEs), presenting a critical challenge to computational efficiency.
Naturally, reutilizing the outcomes of diffusion for successive updates of the 3D object—mimicking the process of traditional 3D reconstruction—could substantially decrease the overall NFEs required. However, traditional 3D reconstruction typically requires a sufficient number of consistently aligned multi-view images, which cannot be directly obtained from the current multi-view diffusion models [24, 47, 58, 52] (see Sec. 6 for a review and discussion). Therefore, similar to existing optimization-based methods, we introduce a reconstruction-based algorithm with multiple rounds of refinement, dubbed score-based iterative reconstruction (SIR), detailed as follows.
Formally, SIR consists of $K$ reconstruction iterations. In the $k$-th iteration, where $k=0,1,\ldots,K-1$, given an initial 3D parameter $\theta_{0}^{(k)}$, we randomly select several camera poses $c^{(k)}$ (detailed in Sec. 4) and employ a rendering function $g(\cdot,\cdot)$ to obtain multi-view images of the current 3D object as follows:
$x^{(k)}=g(\theta^{(k)}_{0},c^{(k)}).$ (7)
For simplicity, let $x^{(k)}$ be a vector by flattening all images and concatenating them together and so do any subsequent set of multi-view images.
As detailed in Sec. 3.2, $x^{(k)}$ is then refined via a pre-trained multi-view diffusion model [47, 24, 58, 52]. The outcome, denoted as $\hat{x}^{(k)}$, serves as the ground truth for the current iteration. In particular, starting from $\theta_{0}^{(k)}$, we optimize the 3D parameters w.r.t. the following reconstruction loss for $I^{(k)}$ steps:
$\mathcal{J}_{\textrm{SIR}}(\theta;c^{(k)},\hat{x}^{(k)})=\|g(\theta,c^{(k)})-\hat{x}^{(k)}\|,$ (8)
[…]
Here we assume the rendered images and those sampled by the diffusion model are the same dimensional for simplicity. We will discuss how to deal with the latent diffusion model (LDM) [43] in Sec. 3.3. Before that, we discuss how to obtain 3D-consistent and high quality $\hat{x}^{(k)}$ from $x^{(k)}$ in Sec. 3.2.
3.2 Obtaining Multi-view Images for Reconstruction by Diffusion
We treat $x^{(k)}$ as noise-free (i.e., at time 00 of the diffusion process) but low-quality images, refined by a forward process followed by a sampling process.
The diffusion model has two equivalent forward processes [51]: the noise-adding process modeled by stochastic differential equations (SDEs) in Eq. (1) and the inversion process based on the probability flow ordinary differential equations (ODEs) in Eq. (4). In comparison, the noise-adding process is more efficient without function evaluation, and it aids in eliminating low-quality details through high-frequency noise. Conversely, the inversion process can preserve the 3D consistency and overall information of the current 3D object. Observations from existing experiments [38] indicate that diffusion models tend to generate details closer to time 00 and overall structure as they approach time $T$. Inspired by this, we design a hybrid forward process that initially adds noise to time $t_{1}^{(k)}$ and then perform DDIM inversion to time $t_{2}^{(k)}$, where $t_{1}^{(k)}\in(0,T)$ and $t_{2}^{(k)}\in[t_{1}^{(k)},T)$ are hyperparameters [^2] Note that if $t^{(k)}_{1}=t^{(k)}_{2}$, the process is purely adding noise.. In contrast to random sampled $t_{2}^{(k)}$ in existing algorithms including SDS [40], we adopt a linearly decreased schedule for $t_{2}^{(k)}$ as $k$ progresses, detailed in Sec. 4. Formally, the process is defined as:
$\tilde{x}^{(k)}=\textrm{Inversion}(\textrm{Noise-adding}(x^{(k)},0\rightarrow t_{1}^{(k)}),t_{1}^{(k)}\rightarrow t_{2}^{(k)}),$ (9)
[…]
Note that generally $x^{(k)}$ does not follow the model distribution defined by the diffusion and the resulting $\tilde{x}^{(k)}$ do not strictly adhere to the marginal distribution at the corresponding time. Nevertheless, we still use existing sampling algorithms to obtain refined images $\hat{x}^{(k)}$ from $\tilde{x}^{(k)}$ as follows:
$\hat{x}^{(k)}={\textrm{Sampler}}(\tilde{x}^{(k)},t_{2}^{(k)}\rightarrow 0).$ (10)
Although other advanced sampling methods [21, 2, 1, 30, 31, 64] exist, we choose the popular DDIM in Eq. (3) and tune its noise hyperparameter $\eta$. The optimal value of $\eta$ depends on the base model and we search for the best one in $\{0,0.5,1\}$ as detailed in Tab. 1. The whole process is presented in Algorithm 1.
Compared with existing methods such as SDS, SIR leverages several NFEs for multiple updates of 3D parameters and lowers the NFEs in total (see Fig. 3(a)).
1
2Input: The number of iterations $K$, an initial 3D object $\theta_{0}^{(0)}$, a set of reconstruction steps $\{I^{(k)}\}_{k=0}^{K-1}$ and a set of camera poses $\{c^{(k)}\}_{k=0}^{K-1}$.
3 Output: A final 3D content $\theta^{(K)}_{0}$. for $k$ from 00 to $K-1$ do
4       Render $N$ images $x^{(k)}=g(\theta^{(k)}_{0},c^{(k)})$ ;
5       Obtain noisy $\tilde{x}^{(k)}$ from forward process for $x^{(k)}$ using Eq. (9);
6       Obtain refined $\hat{x}^{(k)}$ from sampling process for $\tilde{x}^{(k)}$ using Eq. (10);
7       for $i$ from 00 to $I^{(k)}-1$ do
8             Compute loss $L=\|g(\theta,c^{(k)})-\hat{x}^{(k)}\|$ using Eq. (8);
9             Compute the gradient $\nabla L$, update $\theta^{(k)}_{i}$ to $\theta^{(k)}_{i+1}$;
10
11       end for
12      Obtain $\theta^{(k+1)}_{0}=\theta^{(k)}_{I^{(k)}}$.
13 end for
Algorithm 1 Score-based Iterative Reconstruction
3.3 Enabling Optimization in the Pixel Space
Within the widely adopted framework of LDM [43], the rendered images are mapped through an encoder $\mathcal{E}$ [16] to a latent space, where the loss function is calculated. Consequently, the gradients must be propagated back through the encoder, further reducing the efficiency of 3D object generation. In such a case, the SIR loss in the latent space is given by:
$\mathcal{J}_{\textrm{SIR-latent}}(\theta;c^{(k)},\hat{x}^{(k)})$ $=\|\mathcal{E}(g(\theta,c^{(k)}))-\hat{x}^{(k)}\|.$ (11)
An alternative approach maps the diffusion output back to the pixel space via the corresponding decoder $\mathcal{D}$, allowing for a similar loss function to be defined directly in pixel space. This method enables direct updates to 3D parameters without passing gradients through the encoder, thus enhancing efficiency.
However, we argue that this is not feasible for existing optimization methods [40, 56, 60] such as SDS. Even reparameterized as the data predicting form in Eq. (6), the pixel-level quality of the images obtained by SDS is poor for 3D optimization (see Appendix A). In contrast, SIR achieves higher quality generation results through a carefully designed refinement process, thereby enabling optimization in pixel space. The SIR loss in the pixel space is formalized as:
$\mathcal{J}_{\textrm{SIR-pixel}}(\theta;c^{(k)},\hat{x}^{(k)})$ $=\|g(\theta,c^{(k)})-\mathcal{D}(\hat{x}^{(k)})\|,$ (12)
which is about 2-3 times faster than Eq. (11) for 3D parameter optimization, as shown in Fig. 6(a)-6(b). It is set as the default loss throughout the paper.
4 MicroDreamer
In addition to our algorithm SIR, we carefully analyze and design other critical elements in the diffusion model and 3D training. Our comprehensive system is called MicroDreamer to highlight its efficiency.
3D initialization and camera views.
We utilize a direct reconstruction approach to initialize the 3D assets. Specifically, we optimize the loss function in Eq. (8) by several steps (see Tab. 1 for detailed values) to update the 3D content, where $\hat{x}$ represents the images sampled from random noise via the pre-trained multi-view diffusion model. The camera views are uniformly sampled following the corresponding baselines [55, 9] except that the azimuth angles of different views in the same iteration are evenly distributed.
Annealed time schedule.
We utilize an annealed time schedule $\{t_{2}^{(k)}\}_{k=0}^{K-1}$ for the end of the forward process. Intuitively, as the quality of the 3D assets improves, the input $x$ in Eq. (7) becomes more proximate to the model distribution, thus requiring fewer steps to refine. This differs from SDS, which samples uniformly random $t_{2}$. Our preliminary experiments in Fig. 6(e) suggest that a linearly annealed schedule is sufficient. The endpoints of the schedule depend on the diffusion model, detailed in Tab. 1. See Appendix B for the schedule of $t_{1}$.
Reference image loss for image-to-3D.
A reference image is available in image-to-3D, which is regarded as the ground truth front view of the 3D object on Stable Zero123[24, 52] following DreamGaussian [55]. In this way, we add the reference loss in the same form as the reconstruction loss with Eq. (8) in each iteration of training. The weight set for the reference loss can be seen in Tab. 1.
3DGS settings and mesh refinement.
For simplicity, our approach to 3DGS largely adheres to the settings in DreamGaussian[55] unless specified. We incorporate a densify and prune procedure at every 100 updates during the initial 300 updates. We follow the mesh extraction method in DreamGaussian [55] and employ a threshold value of $0.2$ in the Marching Cube algorithm [28, 55]. We utilize SIR to optimize the mesh texture in 3 iterations, with 15 steps per iteration. We use a noise-adding process and DDIM with $\eta=0$ for simplicity.
Table 1: Key hyperparameters of MicroDreamer on three base diffusion models. All models are employed to generate NeRF and the last two are employed to generate 3DGS and mesh. The hyperparameters are shared across the two settings by default. Otherwise, those for 3DGS are shown in brackets. $T$ is end time of diffusion.
Model Select MVDream [47] Stable Zero123 [24, 52] ImageDream [58]
Diffusion
CFG [12] 7.5 3.0 3.0 (2.0)
Forward process hybrid hybrid hybrid (random noise only)
Time schedule of $t_{2}$ $0.8T\rightarrow 0.5T$ $0.8T\rightarrow 0.2T$ $0.8T\rightarrow 0.6T$ ($0.8T\rightarrow 0.4T$)
Sampling process DDIM, $\eta=0.0$ DDIM, $\eta=0.5$ DDIM, $\eta=1.0$
Discretization steps 50 20 10
3D training
Resolution $64\rightarrow 128$ $64\rightarrow 128\rightarrow 196$ (256) $64\rightarrow 128$ (256)
Background learned by NN always white always white
# camera views 4 4 (6) 4
# initialized steps 50 15 50
# iterations $K$ 50 30 (20 or 30) 30
# reconstuction steps $I^{(k)}$ 15 15 15
Loss type $\ell_{1}$ $\ell_{1}$ $\ell_{1}$
Reference color loss - 0.1 (0.3) 0
Reference opacity loss - 0.001 (0.01) 0
5 Experiment
We present the results on NeRF, 3DGS, and ablation sequentially. For a fair comparison, we utilize the widely used CLIP similarity [41] for quantitative comparison following DreamGaussian [55]. We consider 8 views at 0 elevation and evenly distributed in azimuth angles starting from 0. For all methods, the corresponding images are rendered from NeRF in Sec. 5.1 and mesh in Sec. 5.2. We report the average CLIP similarities between these images and the reference image or text. See Appendix C for details of the checkpoint and codebase.
5.1 Results on NeRF
We apply SIR algorithm on NeRF[34, 36, 9] leveraging three multi-view diffusion model from MVDream[47], Stable Zero123[52, 24], and ImageDream[58]. For each model, we selected 6 input prompts from a widely used codebase [9] for testing, calculated NFEs, and recorded the change of the average CLIP similarity in the generation process. Compared with SDS, SIR lowers the total NFEs while generating a 3D object and can accelerate the generation process 5-20 times, as shown in Fig. 3. Besides, SIR can generate NeRF of high visual quality, detailed in Appendix D, and the conclusion agrees with the results in Fig. 3.
(a) Comparsion w.r.t. NFEs.
(b) Comparison on MVDream[47]
(c) Comparison on Stable Zero123[52, 24]
(d) Comparison on ImageDream[58]
Figure 3: Comparison of SIR and SDS on NeRF. (a) shows that SIR lowers the total NFEs. (b-d) show the change of the CLIP similarity in the generation process under different models and SIR achieves a 5-20 times acceleration approximately.
5.2 Results on 3D Gaussian Splatting
Qualitative comparisons.
We show the generated 3D objects amongst various methods and ours in Fig. 4. The baselines include representative feed-forward models One-2-3-45 [23] and TriplaneGaussian [67], and the fastest optimization-based model DreamGaussian [55]. We present the exported mesh from these models. MicroDreamer generates higher quality mesh exported from 3DGS than the feed-forward baselines and halves the time of DreamGaussian [55] while retaining competitive performance.
Figure 4: Qualitative comparisons on mesh from 3DGS. MicroDreamer achieves an excellent efficiency-performance balance in terms of the visual quality of the mesh.
Quantitative comparisons.
We compare MicroDreamer with seven competitive baselines including Point-E [37], Shap-E [14], One-2-3-45 [23], TriplanceGaussian [67], Wonder3D [27], LGM [54] and DreamGaussian [55]. We record the optimization time  [^3] The detailed time recording method is provided in Appendix C. for each model on a single NVIDIA A100 (80GB) GPU, as shown in Fig. 5(a), and compute the average CLIP similarity for mesh on a test dataset consisting of 87 images collected from previous works[23, 25, 55], as shown in Fig. 5(b). As a zero-shot method, MicroDreamer is on par with speed compared to feed-forward methods, faster than DreamGaussian, with a very competitive CLIP similarity. We provide detailed tables in Appendix E.
Notably, the sample quality of MicroDreamer can be improved as the multi-view diffusion evolves. See detailed limitations and future work in Sec. 7.
(a) Optimization time for mesh on a single A100.
(b) CLIP similarity versus inference time.
Figure 5: Quantitative comparisons: efficiency versus quality. MicroDreamer achieves an excellent efficiency-effectiveness balance in terms of the CLIP similarity [41]. The optimization-based methods are highlighted in red in (b).
5.3 Ablation Study
(a) Optimization space [52, 24].
(b) Optimization space [58].
(c) Iterative reconstruction.
(d) Number of camera views.
(e) Time schedule.
(f) Forward process.
Figure 6: Ablation study. (a-b) Optimization in pixel space accelerates the generation process. (c) Iterative reconstruction is necessary. (d) More camera views benefit. (e) A linear schedule is sufficient. (f) A hybrid forward process can be effective.
Ablation on optimization space.
From the results in Fig. 6(a) and Fig. 6(b), we can see that SIR-pixel achieves 2-3 times more acceleration than SIR-latent, verifying the benefits of SIR on enabling optimization in the pixel space in Sec. 3.3.
Ablation on iterative reconstruction.
We highlight the necessity of iterative reconstruction in Fig. 6(c). The case of $K=0$ outputs our 3D initialization, which is reconstructed by a single process from images generated by random noise. Although increasing $K$ improves the quality, we set $K\leq 30$ for efficiency.
Ablation on number of camera poses.
SIR necessitates more than one camera pose for reconstruction, as shown in Fig. 6(d). The performance increases with more camera poses. We choose $6$ on 3DGS with the Stable Zero123 [52] model to balance efficiency and quality. See Tab. 1 for the values in other settings.
Ablation on time schedule.
We compare our linear schedule, the random schedule employed in SDS [40], and a square schedule with $t_{2}^{(k)}=(\frac{k}{K})^{2}\cdot(0.8T-0.2T)+0.2T$. As shown in Fig. 6(e), the schedules perform similarly. For simplicity and efficiency, we use the linear schedule by default.
Ablation on forward process.
Fig. 6(f) compares three forward processes on 3DGS with the Stable Zero123 [52] model. The hybrid strategy performs better than noise-adding and is more efficient than inversion as discussed in Sec. 3.2. See Tab. 1 for other settings.
6 Related work
Optimization-based 3D Generation.
Built upon text-to-image diffusion models, optimization-based approaches [40, 56, 19, 66, 5, 24, 46, 60, 55, 18, 53, 33, 62, 23, 22] generate 3D objects without additional 3D data. Among them, the most relevant line of work [40, 60, 18, 66, 56] proposes various distillation algorithms. Besides SDS [40], ProlificDreamer [60] proposes variational score distillation (VSD) to produce high-fidelity 3D objects via variational inference. LucidDreamer [18] incorporates DDIM inversion to strengthen the forward process. As mentioned in Sec. 1, though promising, such methods suffer from inefficiency problems due to the large number of function evaluations (NFE) and optimization in the latent space. Distinct from existing algorithms, SIR mimics the 3D reconstruction processes to optimize the 3D parameters multiple times given a set of images produced by diffusion and enables optimization in the pixel space. Our experiments demonstrate that SIR is 5-20 times faster than SDS [40] on NeRF and twice faster than the most competitive baseline DreamGaussian [55] on meshes.
Feed-forward 3D Generation.
Contrary to optimization-based methods, feed-forward approaches[3, 4, 6, 14, 37, 26, 13, 35, 59, 61, 65, 67] elect to use large-scale 3D datasets[8, 7] training to achieve a process that can directly generate 3D objects. These methods are characterized by their speed but they often produce lower-quality 3D structures with simple textures. Several recent methods, utilizing more data, have exhibited possibilities. For instance, the large reconstruction model (LRM)[10, 13] demonstrates the feasibility of training a Transformer to achieve a NeRF from a single image. Instant3D[17] further expands on the LRM enabling it to accept multi-view input and reconstruct it into a NeRF. The recent work large multi-view Gaussian model (LGM)[54] employs a similar concept, generating 3DGS from multi-view images. Remarkably, our MicroDreamer is on par with speed compared to such methods trained on extensive 3D data, with a very competitive 3D quality measured by CLIP similarity [41, 55].
Multi-view Prediction based 3D Generation.
There is also a line of work[27, 25] dedicated to enhancing the output of multi-view diffusion models to reconstruct 3D objects. Such systems, like SyncDreamer[25] and Wonder3D[27], train or fine-tune multi-view diffusion models on 3D datasets and use a single reconstruction process to obtain 3D objects from images without iteration. These methods typically rely on 3D prior knowledge and are limited to specific 3D representations [25, 27, 57] with a long reconstruction time. In contrast, thanks to the carefully designed iterative process in SIR, MicroDreamer applies to various 3D representations and is much more efficient (see Fig. 5(a) for a direct comparison).
7 Conclusion
We introduce SIR, an efficient and general algorithm mimicking the 3D reconstruction process to reduce the total NFEs in zero-shot 3D generation. Based on SIR, our method MicroDreamer achieves a 5 to 20 times speed increase for NeRF generation compared to SDS on various base models. Remarkably, MicroDreamer generates high-quality meshes from 3DGS in about 20 seconds, outpacing the fastest zero-shot methods and matching the speed of feed-forward approaches, with a competitive generation quality.
Limitations.
There is potential for further improving the efficiency of MicroDreamer via employing consistency models [50, 32] or alternative sampling models that require fewer steps [20, 63, 45]. Additionally, the fidelity and 3D consistency of the objects produced by MicroDreamer are directly limited by the quality of the outputs from multi-view diffusion. Nevertheless, we believe SIR is promising and may inspire future work as the multi-view diffusion evolves.
Potential Negative Impact.
As a generative approach, the paper can potentially be used for fabricating data and news. Moreover, the 3D content generated from input images may infringe on the privacy or copyright of others. While automatic detection might mitigate these issues, they warrant careful consideration.
8 Acknowledgement
Thanks Cheng Lu for the discussion when he was a PhD student at Tsinghua University.
Appendix 0.A Results of Applying SDS in Pixel Space.
We highlight that SDS optimizing in pixel space can not be effective even if it has the data predicting form by reparameterized as presented in Eq. 6. Fig. 7 shows the poor quality for single-step prediction from the diffusion model using Eq. 2. Therefore, if we attempt to employ SDS for optimization in the pixel space, i.e. optimize 3D objects with 2D images like those generated in Fig. 7, the resulting 3D content would be of subpar quality, as the results presented in Fig. 8.
Figure 7: Results of single-step prediction. For each pair of images, the left is the input image and the right is the output of single-step prediction. The poor quality of the output makes it unsuitable for 3D optimization.
Figure 8: Results of applying SDS in pixel space. SDS optimizing in pixel space can not be effective.
Appendix 0.B Time Schedule of $t_{1}$
The hyperparameter $t_{1}$ regulates the balance between inversion and noise-adding, which varies with different models. In our earlier exploration, we adopted $t_{1}=t^{2}_{2}/T$ for SIR based on MVDream [47] and ImageDream [58] on NeRF, and $t_{1}=0.6t_{2}$ for SIR based on Stable Zero123 [52, 24] on both NeRF and 3DGS. Notably, SIR based on ImageDream on 3DGS opted for $t_{1}=t_{2}$, i.e., adding noise for the forward process. As an advanced multi-view diffusion model, ImageDream can generate images with enhanced 3D consistency, obviating the need for inversion to maintain consistency. Fig. 9 shows some results of SIR for ImageDream on 3DGS using $t_{1}=t_{2}$ schedule.
Figure 9: Results of SIR based on ImageDream on 3DGS. These results images are rendered from 3DGS with no refinement stage.
Appendix 0.C More Implementation Details
We present the codebase, some details in metrics, and the data used in ablation sequentially.
0.C.1 Codebase
For the implementation of SIR and SDS on NeRF, we choose the popularly used framework threestudio [9]. For the implementation of SIR on 3DGS, we follow the framework from DreamGaussian [55] [^4] https://github.com/dreamgaussian/dreamgaussian. The checkpoint we chose for CLIP computation is laion/CLIP-ViT-bigG-14-laion2B-39B-b160k [^5] https://huggingface.co/laion/CLIP-ViT-bigG-14-laion2B-39B-b160k.
In Sec. 5.2, we compare MicroDreamer with seven competitive baselines by utilizing the official code released in GitHub, including Point-E [37] [^6] https://github.com/openai/point-e/, Shap-E [14] [^7] https://github.com/openai/shap-e, One-2-3-45 [23] [^8] https://github.com/One-2-3-45/One-2-3-45, TriplanceGaussian [67] [^9] https://github.com/VAST-AI-Research/TriplaneGaussian, Wonder3D [27] [^10] https://github.com/xxlong0/Wonder3D, LGM [54] [^11] https://github.com/3DTopia/LGM and DreamGaussian [55] [^12] https://github.com/dreamgaussian/dreamgaussian.
0.C.2 Details in Metric
All the results presented in Sec. 5.2 were on the mesh representation. As TriplaneGaussian [67] has no official mesh export code, we apply the mesh exported code from LGM [54] for it.
For extra refine time, we record the time of second stage refinement for MicroDreamer and DreamGaussian [55], the time of generating 3D from multi-view images for Wonder3D[27], and the time for meshing through a NeRF for LGM[54], respectively.
0.C.3 Data Used in Ablation Study
We chose seven images collected from the test dataset to perform our ablation study. We will include these images in the submitted supplementary material.
Appendix 0.D More Results on NeRF
In Fig. 10, we present a comparison of the generated quality between SIR and SDS on NeRF in the same short time under different base diffusion models [47, 52, 24, 58]. SIR can generate NeRF of high visual quality in a shorter time than SDS, as the results represented. Notably, SIR accelerates the generation process in ImageDream more, as it’s an advanced multi-view diffusion model that can produce higher-quality multi-view images and support fewer discretization steps for the diffusion process, lowering NFEs.
Figure 10: Comparison of the generated quality between SIR and SDS in the same short time on NeRF. SIR can generate NeRF of high visual quality in less time than SDS.
Appendix 0.E Quantative Comparison on 3DGS
We provide the detailed quantitative comparison table in Tab. 2. For a fair comparison, we also explored the performance of DreamGaussian with fewer iterations (300 iterations) whose generation time is near MicroDreamer-30 iter. As shown in Tab. 2, MicroDreamer performs better than DreamGaussian over a similar amount of time, further illustrating that MicroDreamer attains a commendable efficiency-performance balance.
Table 2: Quantitative Comparisons. In terms of CLIP similarity, MicroDreamer achieves an excellent balance between efficiency and effectiveness.
Method CLIP-Similarity $\uparrow$ Generation time $\downarrow$ Extra refine $\downarrow$
Point-E 0.510 $\sim 24$s -
Shap-E 0.616 $\sim 5$s -
One-2-3-45 0.617 $\sim 42$s -
Wonder3D 0.699 $\sim 30$s $\sim 140$s
TriplaneGaussian 0.647 $\sim 7$s -
LGM 0.716 $\sim 3$s $\sim 50$s
DreamGaussian 0.698 $\sim 30$s $\sim 10$s
DreamGaussian-300 iter 0.656 $\sim 18$s $\sim 10$s
MicroDreamer-20 iter (Ours) 0.661 $\sim 18$s $\sim 4$s
MicroDreamer-30 iter (Ours) 0.687 $\sim 26$s $\sim 4$s